OpenStudy (anonymous):

Help how do I do this ? How many grams of magnesium metal will react completely with 6.3 liters of 4.5 M HCl? Show all of the work needed to solve this problem. Mg (s) + 2 HCl (aq) yields MgCl2 (aq) + H2 (g)

OpenStudy (cuanchi):

M=n/V n=m/MM stoichimetry of the reaction tells you 2mol HCl/1mol Mg

OpenStudy (jfraser):

Use the concentration and volume of acid to find the moles of acid available, as @cuanchi says. Then use the balanced reaction to find out the moles of Mg are needed. Then use the molar mass of Mg to find the mass of Mg that is needed